2 years before father was twice as old as his son and hence after 2 years thrice the age of father equals to 5 times the age of son.find their present ages?
Answer by ankor@dixie-net.com(22740)

2 years before father was twice as old as his son and hence after 2 years thrice the age of father equals to 5 times the age of son.find their present ages?
:
Let f = father's present age
Let s = son's present age
:
Write an equation for each statement:
:
" 2 years before father was twice as old as his son"
f - 2 = 2(s - 2)
f - 2 = 2s - 4
f = 2s - 4 + 2
f = 2s - 2
:
"and hence after 2 years thrice the age of father equals to 5 times the age of son."
3(f+2) = 5(s+2)
3f + 6 = 5s + 10
3f = 5s + 10 - 6
3f = 5s + 4
Replace f with (2s-2)
3(2s-2) = 5s + 4
6s - 6 = 5s + 4
6s - 5s = 4 + 6
s = 10 is the son's present age
:
I'll let you find f, check the solutions in the given statements